# Atomic-Scale Mechanisms of Li-Ion Transport Mediated by Li10GeP2S12 in Composite Solid Polyethylene Oxide Electrolytes

### v1

@@ Abstract
Polymer electrolytes incorporating Li10GeP2S12 (LGPS) nanoparticles are promising for solid-state lithium batteries due to their potential for enhanced ionic conductivity; yet, the atomistic mechanisms driving this enhancement remain debated. Here, we systematically investigate the relationship between LGPS nanoparticle loading, polyethylene oxide microstructure, and Li-ion transport using a combination of molecular dynamics (MD) simulations, experimental ionic conductivity measurements, and density functional theory (DFT) calculations. MD simulations and experiments reveal good agreement on ionic conductivity as a function of LGPS concentrations of up to 10 weight % (x%), exhibiting a volcano-like curve with ionic conductivity increasing fivefold from the low concentrations and can be accounted for by a classical transport mechanism governed by polymer segmental dynamics and interface effects. However, at more than 10% LGPS, experiments show further conductivity enhancement that cannot be accounted for by MD simulations, indicating a shift to another transport mechanism. DFT calculations elucidate that, at the polymer|LGPS interface, Li-ion migration proceeds via vacancy-driven hopping, with barriers sensitive to local atomic composition-low-barrier pathways are possible when S atoms dominantly occupy the sites on the interface to facilitate Li-hopping, while pathways involving Ge act as obstacles to Li transport. These results establish that optimized interfacial chemistry and electrolyte structure enable efficient, barrier-lowering migration channels that are distinct from bulk polymer or ceramic behavior. Our approach reconciles experiments with classical simulations at low LGPS concentrations and quantum chemical interface calculations, highlighting design criteria for maximizing the performance of these types of solid composite polymer electrolytes and guiding the development of advanced lithium batteries.
1 Introduction
@@
All-solid-state batteries (ASSBs) are among the most promising next-generation energy storage systems, offering high energy density, enhanced safety, and long cycle life [kondori2023room, zhang2025fiber, zhou2025molecular, lemaalem2023graphyne, lemaalem2023tunable]. Within ASSBs technology, solid composite polymer electrolytes (CPEs), which incorporate inorganic lithium-ion conductors into polymer matrices, have attracted significant attention for their ability to combine the mechanical flexibility of polymers with the high ionic conductivity of inorganic phases. Among the various inorganic fillers, Li10GeP2S12 (LGPS) stands out as one of the most conductive sulfide electrolytes, exhibiting room temperature ionic conductivities exceeding $10^{-2}\,\text{S\,cm}^{-1}$ [Kato2016, Seino2014, Kuhn2013, Yu2021, Schweiger2022].
@@
A wide range of experimental and theoretical studies have explored the ion transport mechanisms in solid polymer composite electrolytes containing nanoparticles. The addition of inorganic fillers such as LGPS can disrupt polymer crystallinity, enhance the amorphous content, and increase segmental motion, thereby facilitating cation migration [Zhou2018, Bhattacharyya2020, zhang2019sulfide]. At moderate loadings, nanoparticles may create percolative ion-conduction networks and provide high-dielectric environments that increase the concentration of free lithium ions [Dawson2022, Cui2023, zhang2019sulfide]. Several reports also highlight that the interfacial region between the polymer and nanoparticles is critical: ion transport can be improved through tailored nanoparticle-polymer interactions, as well as functionalization of the filler to optimize compatibility and interfacial ionic conduction pathways [zhang2019sulfide, Hori2022, pan2020flexible]. However, excessive nanoparticle content often leads to agglomeration, restricted polymer dynamics, and diminished conductivity, as the nanoparticles can immobilize nearby polymer chains and hinder segmental motion [Zhou2018, Bhattacharyya2020]. Morphological control of the filler, such as using nanowires or aligned frameworks rather than particles, has also been shown to provide continuous and highly conductive pathways for lithium ions [zhang2019sulfide, Dawson2022].
@@
Nevertheless, the precise atomistic mechanisms underpinning ion transport in nanoparticle-polymer composite electrolytes, including the role of the interface and optimal filler concentration, remain insufficiently understood. In this work, we employ a combined atomistic molecular dynamics (MD) and density functional theory (DFT) framework to elucidate the multiscale mechanisms governing ion transport in a CPE comprising LGPS nanoparticles, polyethylene oxide (PEO), methoxy-poly(ethylene oxide)-trimethoxysilane (mPEO-TMS), and Lithiumbis(trifluoromethanesulfonyl)imide (LiTFSI) used in a solid-state Li-air battery [kondori2023room]. The mPEO-TMS is used to functionalize the LGPS nanoparticles to stabilize them and enhance interfacial conduction. This CPE is similar to another one developed for use in a Li-ion battery [pan2020flexible]. In that study, the LGPS functionalization was performed using polyethylene glycol (PEG) and (3-chloropropyl)trimethoxysilane (CTMS).
@@ 2.1 MD Simulation Details
Atomistic molecular dynamics (MD) simulations were performed to model a composite solid polymer electrolyte comprising Li10GeP2S12 (LGPS) nanoparticles, poly(ethylene oxide) (PEO), methoxy-poly(ethylene oxide)-trimethoxysilane (mPEO-TMS), and lithium bis(trifluoromethanesulfonyl)imide (LiTFSI) (Figure 1), based on the experimental composition presented in the supplementary information (Tables S1 and S2). In this study, we investigated a broader range of LGPS loadings from 0 to 40 weight % (x%) (relative to the total weight of all electrolyte components, except the salt) to assess the effect of filler content on ion transport. For computational feasibility, the LGPS particle size of 1.2 nm was used while preserving the experimental mass ratios among all components. We note that the experimental LGPS crystallite size is 17 nm, but the actual nanoparticle size is not known [kondori2023room]. When rescaling the LGPS nanoparticle size, the molecular weight of the PEO chains was also proportionally modified to match the LGPS-rescaled molecular weight, ensuring consistent polymer–filler mass relationships. The other small-molecule species, namely mPEO-TMS and LiTFSI, retained their original molecular models; only their quantities were adjusted to ensure that the experimental weight composition was strictly maintained. The simulation cell contained a set number of LGPS nanoparticles dispersed in polymer-salt matrix, as presented in Table 1. All simulations were carried out using LAMMPS [thompson2022lammps] with the OPLS-AA force field for organic species [doherty2017revisiting] and the UFF potential for LGPS and TMS [rappe1992uff], more computational details are discussed in the Supplementary Information.
@@ 2.2 DFT calculations details
We carried out spin-polarized DFT calculations to optimize the structures as well as to determine the Li cation migration pathway. The initial LGPS bulk structure was sourced from the Materials Project and cut using Crystal Maker to a 1.2 nm particle comprising 400 atoms [Jain2013]. The DFT calculations implemented the Projector-Augmented Wave (PAW) potentials for the elements as supplied by the Vienna Ab Initio Simulation Package (VASP) [Kresse1996b, Kresse1999]. We used the Generalized Gradient Approximation (GGA), based on the Perdew–Burke–Ernzerhof approach, to account for the exchange-correlation energies [Perdew1996]. The kinetic energy cutoff was set to 600 eV to enhance accuracy, and the Brillouin zone was sampled at the gamma point with the k-mesh set based on the respective lattice parameters. The energy convergence thresholds for the electronic self-consistent loop and ionic relaxation loop were set to 0.01 meV/Å and 0.1 meV/Å, respectively. All the structures were visualized using VESTA [Momma2011].
@@
The surface of the LGPS nanoparticles was modified using the silane coupling agent mPEO-TMS. In Ref. Ref [kondori2023room] evidence was presented that the Si atoms in mPEO-TMS form bonds with the S atoms in LGPS, thereby stabilizing the nanoparticle to prevent decomposition at the Li anode or cathode [kondori2023room]. As illustrated in Figure 2, this surface reaction results in the formation of SiS3 bonds and the release of three LiOCH3 molecules. The mPEO-TMS/LGPS structure based on a 1.2 nm particle, with its geometry optimized, was used to calculate energy barriers for the Li cation transport along the mPEO-TMS|LGPS interface. We used the nudged elastic band (NEB) method described by Sheppard et al. [Sheppard2008] to determine the minimum energy paths. The NEB method uses a string of images (geometric configurations with varying atomic positions) between the initial and final states to describe the migration pathway. These images are connected by spring forces, which ensure equal spacing along the path. They are minimized using DFT and compared with the initial state energy to determine the energy required for the ion to move along the interface.
@@ 2.3 Ionic conductivity measurements
The ionic conductivity of the CPE was measured by electrochemical impedance spectroscopy (EIS, AC impedance) using a symmetric SS||CPE||$S$ cell, following the procedure described in Ref. [kondori2023room]. Measurements were performed at room temperature over a frequency range of 100 kHz to 0.1 Hz with an AC amplitude of 5 mV using a BioLogic SP150 potentiostat, which provides an adequate signal-to-noise ratio while maintaining a linear current response. The ionic conductivity $\sigma$ of the CPE was measured as: $\sigma=\frac{L}{A\,R_{\mathrm{s}}}$, where $L$ (cm) and $A$ (cm2) are the thickness and the electrode–electrolyte contact area of the CPE, respectively, and $R_{\mathrm{s}}$ ($\Omega$) is the bulk resistance. Ionic conductivity from these measurements are shown in Figure 3 for the solid electrolyte composed of poly(ethylene oxide) (PEO), 3-[methoxy(polyethyleneoxy)6-9]propyltrimethoxysilane (mPEO), and Li10GeP2S12 (LGPS) as a function of LGPS content. The LGPS mass was varied from 0 mg to 400 mg, while the mPEO mass was kept constant at 1 g. In all samples, 0.5 g of PEO with a molecular weight of 1M was used. These measurements were also carried out in an SS||CPE||$S$ cell at room temperature.
@@ 3.1 Transport properties from MD simulation
Figure 3: Ionic conductivity as a function of the LGPS weight ratio, with respect to the experimental composition presented in Table S2 and scaled for computational feasibility in Table 1, from MD simulations compared with experimental data (see Figure S1 for the experimental data shown in this plot as well as other data). Note that the result for 0 wt% is from Ref. [kondori2023room] and the 3.2% results differs slightly from that in Ref. [kondori2023room] due to uncertainties in measurements (see SI).
@@
Here, $q$ is the elementary charge, $z_{i}$ is the charge number (valence) of ion $i$, $\mathbf{v}_{i}$ is the velocity of ion $i$, $T$ is the temperature, $k_{B}$ is the Boltzmann constant, $V$ is the volume of the simulation box, and $N$ is the number of ions. The Green-Kubo relations can also be expressed in terms of particle positions rather than velocities. This form, analogous to computing self-diffusion coefficients from the mean-squared displacement of particle positions, is used to compute $L^{ij}$ in this work [lemaalem2023effects]:
@@
The molecular dynamics (MD) simulations for varying LGPS nanoparticle concentrations reveal a complex dependence of ionic conductivity on LGPS nanoparticle loading (see Figure 3). Experimental data for a similar system based on LGPS particles, LiTFSI, mPEO-TMS, and PEO is shown in Figure 3. In this figure, the measured ionic conductivities are plotted as a function of LGPS loading for comparison with the calculated values. Details of the experimental measurements are given in the SI. For LGPS weight fractions up to 10 x%, the ionic conductivities calculated from MD simulations using the Green-Kubo relation agree with the measured values shown in (Figure 3), demonstrating a sharp fivefold increase in conductivity between 0 and 3.2 x% LGPS, followed by a decline past this optimal loading (Figure 3). This behavior reflects classical conduction mechanisms, enhanced segmental mobility, nanoconfinement effects, and favorable polymer/nanoparticle interfaces, that promote ion transport and are well-captured by MD. Polymer chain and ionic diffusion coefficients are highly correlated (Figure 4(a)), suggesting a global enhancement of polymer dynamics due to LGPS addition. The potential effect of mPEO-TMS inclusion on making the LGPS surface more amenable to ion transport is shown in the DFT calculations reported in Section 3.3. The MD simulations show that Li+ ions diffuse more rapidly than TFSI- (Figure 4(b)), indicating the selective formation of conduction pathways along polymer and nanoparticle interfaces. These results are consistent with the literature: LGPS nanoparticles improve conductivity via interfacial interactions that increase free volume, reduce crystallinity, and optimize Lewis acid–base effects [fu2022ion, sand2024critical, zhenginterface].
@@
The two ionic conduction regimes deduced from the MD simulations (Figure 1) and from the experiments (Figure 3) are illustrated in Figure 5. In regime I, the conduction pathways are through the polymers and the polymer/LGPS interfaces, giving rise to a volcano type curve as a function of LGPS wt%. The MD simulations suggest that the decrease in ionic conductivity above 3.2 wt% results from the formation of LGPS clusters that interrupt polymer-based pathways, whereas the increase below 3.2 wt% reflects the enhanced contribution of polymer/interface transport and segmental diffusion [rajahmundry2024understanding, choo2020diffusion], which is a polymer-type diffusion process associated with the local motion of polymer chain segments (Figure 4). At LGPS contents greater than 20 wt% the experimental results show that the ionic conductivity starts increasing again. The MD simulations, however, do not reproduce this experimentally observed increase at LGPS nanoparticle loadings above 20 wt% (Figure 3). In this high-loading regime, ionic transport likely occurs primarily through LGPS agglomerates and/or between percolating LGPS nanoparticle clusters, a mechanism not captured in the current MD model, as schematically illustrated in regime II in Figure 5.

### v2

@@ Abstract
Polymer electrolytes incorporating Li10GeP2S12 (LGPS) nanoparticles are promising for solid-state lithium batteries due to their potential for enhanced ionic conductivity; yet, the atomistic mechanisms driving this enhancement remain debated. Here, we systematically investigate the relationship between LGPS nanoparticle loading, polyethylene oxide microstructure, and Li-ion transport using a combination of molecular dynamics (MD) simulations, experimental ionic conductivity measurements, and density functional theory (DFT) calculations. MD simulations and experiments reveal good agreement on ionic conductivity as a function of LGPS concentrations of up to 10 weight % (x%), exhibiting a volcano-like curve with ionic conductivity increasing fivefold from the low concentrations and can be accounted for by a classical transport mechanism governed by polymer segmental dynamics and interface effects. However, at more than 10% LGPS, experiments show further conductivity enhancement that cannot be accounted for by MD simulations, indicating a shift to another transport mechanism. DFT calculations elucidate that, at the polymer|LGPS interface, Li-ion migration proceeds via vacancy-driven hopping, with barriers sensitive to local atomic composition-low-barrier pathways are possible when S atoms dominantly occupy the sites on the interface to facilitate Li-hopping, while pathways involving Ge act as obstacles to Li transport. These results establish that optimized interfacial chemistry and electrolyte structure enable efficient, barrier-lowering migration channels that are distinct from bulk polymer or ceramic behavior. Our approach reconciles experiments with classical simulations at low LGPS concentrations and quantum chemical interface calculations, highlighting design criteria for maximizing the performance of these types of solid composite polymer electrolytes and guiding the development of advanced lithium batteries. {tocentry} [Uncaptioned image]
Introduction
@@
All-solid-state batteries (ASSBs) are among the most promising next-generation energy storage systems, offering high energy density, enhanced safety, and long cycle life 10, 31, 33, 16, 15. Within ASSBs technology, solid composite polymer electrolytes (CPEs), which incorporate inorganic lithium-ion conductors into polymer matrices, have attracted significant attention for their ability to combine the mechanical flexibility of polymers with the high ionic conductivity of inorganic phases. Among the various inorganic fillers, Li10GeP2S12 (LGPS) stands out as one of the most conductive sulfide electrolytes, exhibiting room temperature ionic conductivities exceeding $10^{-2}\,\text{S\,cm}^{-1}$ 9, 25, 13, 29, 24.
@@
A wide range of experimental and theoretical studies have explored the ion transport mechanisms in solid polymer composite electrolytes containing nanoparticles. The addition of inorganic fillers such as LGPS can disrupt polymer crystallinity, enhance the amorphous content, and increase segmental motion, thereby facilitating cation migration 34, 1, 30. At moderate loadings, nanoparticles may create percolative ion-conduction networks and provide high-dielectric environments that increase the concentration of free lithium ions 4, 3, 30. Several reports also highlight that the interfacial region between the polymer and nanoparticles is critical: ion transport can be improved through tailored nanoparticle-polymer interactions, as well as functionalization of the filler to optimize compatibility and interfacial ionic conduction pathways 30, 7, 19. However, excessive nanoparticle content often leads to agglomeration, restricted polymer dynamics, and diminished conductivity, as the nanoparticles can immobilize nearby polymer chains and hinder segmental motion 34, 1. Morphological control of the filler, such as using nanowires or aligned frameworks rather than particles, has also been shown to provide continuous and highly conductive pathways for lithium ions 30, 4.
@@
Nevertheless, the precise atomistic mechanisms underpinning ion transport in nanoparticle-polymer composite electrolytes, including the role of the interface and optimal filler concentration, remain insufficiently understood. In this work, we employ a combined atomistic molecular dynamics (MD) and density functional theory (DFT) framework to elucidate the multiscale mechanisms governing ion transport in a CPE comprising LGPS nanoparticles, polyethylene oxide (PEO), methoxy-poly(ethylene oxide)-trimethoxysilane (mPEO-TMS), and Lithiumbis(trifluoromethanesulfonyl)imide (LiTFSI) used in a solid-state Li-air battery 10. The mPEO-TMS is used to functionalize the LGPS nanoparticles to stabilize them and enhance interfacial conduction. This CPE is similar to another one developed for use in a Li-ion battery 19. In that study, the LGPS functionalization was performed using polyethylene glycol (PEG) and (3-chloropropyl)trimethoxysilane (CTMS).
@@ 1.1 MD Simulation Details
Atomistic molecular dynamics (MD) simulations were performed to model a composite solid polymer electrolyte comprising Li10GeP2S12 (LGPS) nanoparticles, poly(ethylene oxide) (PEO), methoxy-poly(ethylene oxide)-trimethoxysilane (mPEO-TMS), and LiTFSI salt (Figure 1), based on the experimental composition presented in the supplementary information (Tables S1 and S2). In this study, we investigated a broader range of LGPS loadings from 0 to 40 weight % (x%) (relative to the total weight of all electrolyte components, except the salt) to assess the effect of filler content on ion transport. For computational feasibility, the LGPS particle size of 1.2 nm was used while preserving the experimental mass ratios among all components. We note that the experimental LGPS crystallite size is 17 nm, but the actual nanoparticle size is not known 10. When rescaling the LGPS nanoparticle size, the molecular weight of the PEO chains was also proportionally modified to match the LGPS-rescaled molecular weight, ensuring consistent polymer–filler mass relationships. The other small-molecule species, namely mPEO-TMS and LiTFSI, retained their original molecular models; only their quantities were adjusted to ensure that the experimental weight composition was strictly maintained. The simulation cell contained a set number of LGPS nanoparticles dispersed in polymer-salt matrix, as presented in Table 1. All simulations were carried out using LAMMPS 27 with the OPLS-AA force field for organic species 5 and the UFF potential for LGPS and TMS 22, more computational details are discussed in the Supplementary Information.
@@ 1.2 DFT calculations details
We carried out spin-polarized DFT calculations to optimize the structures as well as to determine the Li cation migration pathway. The initial LGPS bulk structure was sourced from the Materials Project and cut using Crystal Maker to a 1.2 nm particle comprising 400 atoms 8. The DFT calculations implemented the Projector-Augmented Wave (PAW) potentials for the elements as supplied by the Vienna Ab Initio Simulation Package (VASP) 11, 12. We used the Generalized Gradient Approximation (GGA), based on the Perdew–Burke–Ernzerhof approach, to account for the exchange-correlation energies 20. The kinetic energy cutoff was set to 600 eV to enhance accuracy, and the Brillouin zone was sampled at the gamma point with the k-mesh set based on the respective lattice parameters. The energy convergence thresholds for the electronic self-consistent loop and ionic relaxation loop were set to 0.01 meV/Å and 0.1 meV/Å, respectively. All the structures were visualized using VESTA 18.
@@
The surface of the LGPS nanoparticles was modified using the silane coupling agent mPEO-TMS. In Ref. 10 evidence was presented that the Si atoms in mPEO-TMS form bonds with the S atoms in LGPS, thereby stabilizing the nanoparticle to prevent decomposition at the Li anode or cathode 10. As illustrated in Figure 2, this surface reaction results in the formation of SiS3 bonds and the release of three LiOCH3 molecules. The mPEO-TMS/LGPS structure based on a 1.2 nm particle, with its geometry optimized, was used to calculate energy barriers for the Li cation transport along the mPEO-TMS|LGPS interface. We used the nudged elastic band (NEB) method described by Sheppard et al. 26 to determine the minimum energy paths. The NEB method uses a string of images (geometric configurations with varying atomic positions) between the initial and final states to describe the migration pathway. These images are connected by spring forces, which ensure equal spacing along the path. They are minimized using DFT and compared with the initial state energy to determine the energy required for the ion to move along the interface.
@@ 1.3 Ionic conductivity measurements
The ionic conductivity of the CPE was measured by electrochemical impedance spectroscopy (EIS, AC impedance) using a symmetric SS||CPE||$S$ cell, following the procedure described in Ref. 10. Measurements were performed at room temperature over a frequency range of 100 kHz to 0.1 Hz with an AC amplitude of 5 mV using a BioLogic SP150 potentiostat, which provides an adequate signal-to-noise ratio while maintaining a linear current response. The ionic conductivity $\sigma$ of the CPE was measured as: $\sigma=\frac{L}{A\,R_{\mathrm{s}}}$, where $L$ (cm) and $A$ (cm2) are the thickness and the electrode–electrolyte contact area of the CPE, respectively, and $R_{\mathrm{s}}$ ($\Omega$) is the bulk resistance. Ionic conductivity from these measurements are shown in Figure 3 for the solid electrolyte composed of poly(ethylene oxide) (PEO), 3-[methoxy(polyethyleneoxy)6-9]propyltrimethoxysilane (mPEO), and Li10GeP2S12 (LGPS) as a function of LGPS content. The LGPS mass was varied from 0 mg to 400 mg, while the mPEO mass was kept constant at 1 g. In all samples, 0.5 g of PEO with a molecular weight of 1M was used. These measurements were also carried out in an SS||CPE||$S$ cell at room temperature.
@@ 2.1 Transport properties from MD simulation
Figure 3: Ionic conductivity as a function of the LGPS weight ratio, with respect to the experimental composition presented in Table S2 and scaled for computational feasibility in Table 1, from MD simulations compared with experimental data (see Figure S1 for the experimental data shown in this plot as well as other data). Note that the result for 0 wt% is from Ref. 10 and the 3.2% results differs slightly from that in Ref. 10 due to uncertainties in measurements (see SI).
@@
Here, $q$ is the elementary charge, $z_{i}$ is the charge number (valence) of ion $i$, $\mathbf{v}_{i}$ is the velocity of ion $i$, $T$ is the temperature, $k_{B}$ is the Boltzmann constant, $V$ is the volume of the simulation box, and $N$ is the number of ions. The Green-Kubo relations can also be expressed in terms of particle positions rather than velocities. This form, analogous to computing self-diffusion coefficients from the mean-squared displacement of particle positions, is used to compute $L^{ij}$ in this work 14:
@@
The molecular dynamics (MD) simulations for varying LGPS nanoparticle concentrations reveal a complex dependence of ionic conductivity on LGPS nanoparticle loading (see Figure 3). Experimental data for a similar system based on LGPS particles, LiTFSI, mPEO-TMS, and PEO is shown in Figure 3. In this figure, the measured ionic conductivities are plotted as a function of LGPS loading for comparison with the calculated values. Details of the experimental measurements are given in the SI. For LGPS weight fractions up to 10 x%, the ionic conductivities calculated from MD simulations using the Green-Kubo relation agree with the measured values shown in (Figure 3), demonstrating a sharp fivefold increase in conductivity between 0 and 3.2 x% LGPS, followed by a decline past this optimal loading (Figure 3). This behavior reflects classical conduction mechanisms, enhanced segmental mobility, nanoconfinement effects, and favorable polymer/nanoparticle interfaces, that promote ion transport and are well-captured by MD. For computational feasibility, the LGPS particle size was rescaled to 1.2 nm. We acknowledge that this results in a higher surface-to-volume ratio compared to experimental crystallites ($\approx$ 17 nm), which may quantitatively amplify interfacial ion dynamics and the resulting cation transference number ($t_{+}$). However, this scaling preserves the qualitative physics of polymer segmental dynamics and interfacial saturation essential for understanding the low-loading regime. Polymer chain and ionic diffusion coefficients are highly correlated (Figure 4(a)), suggesting a global enhancement of polymer dynamics due to LGPS addition. We note that the excess mPEO-TMS, which exceeds the amount required for surface passivation, acts as an active component of the bulk electrolyte. Its structural similarity to PEO ensures miscibility, contributing to the bulk conductivity in the polymer-rich regime. The potential effect of mPEO-TMS inclusion on making the LGPS surface more amenable to ion transport is shown in the DFT calculations reported in Section 2.3. The MD simulations show that Li+ ions diffuse more rapidly than TFSI- (Figure 4(a) and (b)), indicating the selective formation of conduction pathways along polymer and nanoparticle interfaces. These results are consistent with the literature: LGPS nanoparticles improve conductivity via interfacial interactions that increase free volume, reduce crystallinity, and optimize Lewis acid–base effects 6, 23, 32.
@@
The two ionic conduction regimes deduced from the MD simulations (Figure 1) and from the experiments (Figure 3) are illustrated in Figure 5. In regime I, the conduction pathways are through the polymers and the polymer/LGPS interfaces, giving rise to a volcano type curve as a function of LGPS wt%. The MD simulations suggest that the decrease in ionic conductivity above 3.2 wt% results from the formation of LGPS clusters that interrupt polymer-based pathways, whereas the increase below 3.2 wt% reflects the enhanced contribution of polymer/interface transport and segmental diffusion 21, 2, which is a polymer-type diffusion process associated with the local motion of polymer chain segments (Figure 4). At LGPS contents greater than 20 wt% the experimental results show that the ionic conductivity starts increasing again. The discrepancy between MD simulations and experimental results at loadings above 20% wt highlights the limitations of classical molecular dynamics. The classical force fields (UFF/OPLS-AA) treat LGPS particles as high-impedance materials because they lack appropriate parameterization for ion diffusion through the LGPS bulk and for the vacancy-mediated Li hopping mechanism at the LGPS surface. Consequently, classical MD fails to capture the experimentally observed recovery in ionic conductivity at high loadings ( $\geq$ 20% wt), where transport becomes dominated by the inorganic phase. This regime involves Li-ion conduction through percolated ceramic networks, consistent with the spectral signatures reported by Li et al. 17. We note that the present DFT calculations describe only the surface diffusion mechanism and do not account for bulk ion transport.